Can someone please help me with this problem?
A man sold 2 acres more than 3/5 of his farm and had 4 acres less than 1/2 of it left. How many acres in the farm? Call this X. Find X.

Answer by Alan3354(69443)

Ok, it's x.
x - ( (3/5)x + 2) = (1/2)x - 4
x - 3x/5 - 2 = x/2 - 4
2x/5 - 2 = x/2 - 4
4x - 20 = 5x - 40
x = 20
------